OpenStudy (anonymous):

Given [h(x)=x^5 -2, g(x)=x^4 -1, f(x)=x^2 -2 find f(g(1)+h(1))

OpenStudy (anonymous):

g(1) = (1^4)-1 = 1-1 = 0 h(1) = (1^5)-2 = 1-2 = -1 f(-1+0) = f(-1) f(-1) = (-1)^2-2 = 1-2 = -1

OpenStudy (anonymous):

thanls